OpenStudy (anonymous):

The diagonal of a square tile is 4 inches longer than its side. Find the length of the side.

OpenStudy (anonymous):

using the pythagoran theorem \[16s ^{2} = \sqrt{\left( 2s ^{2} \right)}\] where s is the side of the square solving for s we get \[s =\sqrt{2}\div16\]

OpenStudy (anonymous):

thanks!